OpenStudy (sophadof):

6^1/2 in Radical Form Show steps

OpenStudy (anonymous):

sqrt of 6

TheSmartOne (thesmartone):

Hint: \(\sf\Huge \sqrt[x]{y^z} = y^{\frac{z}{x}}\)

OpenStudy (anonymous):

i would love to travel 9.4 million times the speed of light